OpenStudy (anonymous):

If f(x) = 7x - 1, what does f(12) represent? @jordanloveangel

OpenStudy (anonymous):

substitute 12 fox x and solve f(x)=7*12-1=84-1=83

OpenStudy (anonymous):

can you afn me?

TheSmartOne (thesmartone):

So f(x)=7x-1 Bascially everywhere you see an x, you just substitute 12 So we get f(12)=7(12)-1 And you can solve 7*12-1 What ever you get from 7*12-1 will be your answer.